# RoTBench: A Multi-Level Benchmark for Evaluating the Robustness of Large Language Models in Tool Learning

### v2

@@ Abstract
RoTBench: A Multi-Level Benchmark for Evaluating the Robustness
of Large Language Models in Tool Learning
[…]
@@ 1 Introduction
Tool learning has emerged as a critical concept for empowering large language models (LLMs) Brown et al. (2020); Bai et al. (2022); Touvron et al. (2023a) to interact with the real world Yang et al. (2023); Mialon et al. (2023); Qin et al. (2023a). In this context, the external environment of an LLM contains an ensemble of integrated tools. Each tool is uniquely identified by its name and is described by a succinct paragraph that explains its functionality. Similarly, every parameter within these tools is characterized by its name, along with a description that clarifies its purpose, its optionality, and other pertinent details. For illustration, Figure 1 represents information of three distinct tools.
Figure 1: Information of three tools. The tools are used for executing code, querying the weather, and searching for information, where “code” and “te” are required parameters and “location” is optional.
Recent research has centered on examining how well LLMs can effectively employ tools within a carefully designed and stable environment. From one perspective, specific studies have scrutinized the outcomes of LLMs’ tool usage, verifying both the accuracy of tool selection and the efficacy of the generated responses Qin et al. (2023b); Huang et al. (2023). This analysis involved evaluating the relevance of the selected tools and the final responses in fulfilling users’ requirements. On the other hand, other investigations have delved into the intricate process of tool utilization by LLMs, striving for a more comprehensive assessment of their performance in tool learning Chen et al. (2023d); Ye et al. (2024). This includes an analysis of the diverse capabilities necessary for LLMs to excel in tool learning while also identifying any limitations they may have in this regard.
However, these studies fail to account for the robustness of LLMs in the face of inevitable noise in real-world scenarios Chen et al. (2023b); Liu et al. (2023). Consider the tool labeled “ABC” in Figure 1, despite its name lacking specific meaning, it remains capable of effectively serving weather information queries. Similarly, in the “search” tool, the parameter “te” is assigned the role of a user query and becomes a mandatory parameter. Consequently, it becomes imperative to investigate whether LLMs can proficiently identify these tools and skillfully configure parameters to meet user needs in noisy real-world environments. This research is essential to guarantee their reliability in practical applications.
@@
2 RoTBench
As depicted in Figure 2, RoTBench encompasses five environments, each characterized by varying levels of noise, facilitating a thorough evaluation of the robustness of LLMs throughout the three stages of tool usage. Detailed descriptions are provided in the following sections.
2.1 Data Collection
@@
Table 1: Statistics information of the collected data. “# Sce”, “# Query”, “# Cat”, “# Subcat”, and “# Tool” correspond to the count of scenarios, user queries, tool categories, tool subcategories, and individual tools, respectively.
In order to thoroughly cater to real-world requirements and encompass commonly utilized tools, we utilize ToolEyes Ye et al. (2024), an evaluation system designed for tool learning. This system defines seven real-world application scenarios that span text generation, data understanding, real-time search, application manipulation, personal life, information retrieval, and financial transactions.
Within each of these scenarios, we have randomly selected 15 user requirements for analysis. Since the raw data offers tool information without standardized invocation paths, we have manually labeled these paths to facilitate the evaluation process. Detailed statistics of the data can be found in Table 1.
@@ 2.2 Environments Construction
In order to thoroughly evaluate the robustness of LLMs in tool learning, we create five external environments with varying levels of noise: Clean, Slight, Medium, Heavy, and Union. These noise apply to both the tool and its parameters.
Clean
@@
Slight
Slight-level environment encompasses three types of noise: insertion, omission, and substitution. These correspond to real-world occurrences such as an excess of characters, missing characters, and character errors when naming tools or parameters. To minimize the impact of this noise on the overall environmental information, we introduce noise in the following ways: 1) We randomly select half of the available tools within the environment. For these selected tools, a random form of noise is applied, altering up to 1/3 of the characters. The remaining tools remain unaltered, resulting in the creation of 105 new data points. 2) For each tool, we randomly select half of the parameters and introduce noise into their names using the method described above. The remaining parameters remain unchanged, generating an additional 105 new data entries. By combining these two approaches, we create a Slight-level environmental test set consisting of 210 test cases.
Medium
Medium-level environment introduces two types of noise: reversal and nonsense. These mirror real-world scenarios where names are reversed or replaced with random strings, rendering the information meaningless within the environment. To apply noise, we follow these procedures: 1) We randomly select half of the available tools. For these tools, there is a 50% probability that their names will be substituted with random strings, each containing up to 10 characters. Additionally, there is a 50% chance that the names of these tools will be reversed. This process yields 105 test cases. 2) For each tool, half of the parameters are randomly chosen. These parameters may undergo a 50% chance of having their names substituted with random strings, each containing up to 5 characters, or a 50% chance of being reversed. This leads to 105 test cases. It is worth noting that if the reversal process does not alter the name, it will be replaced with a random string. Consequently, we have successfully generated 210 test cases for the Medium-level environment.
Heavy
Union
2.3 Staged Evaluation
The process of LLMs’ tool learning contains three crucial stages: tool selection, parameter identification, and content filling. We evaluate the robustness performance of LLMs at each of these stages and analyze their respective variations.
Tool Selection
$s_{TS}=\mathbb{I}(t=\hat{t})$ (1)
Parameter Identification
Parameter identification involves recognizing the required parameters and outputting their respective names based on their specified needs, following the selection of the appropriate tool. This process necessitates choosing the mandatory parameters, while the optional ones are selected based on actual requirements. Similar to tool selection, the name of the parameter serves as an identifier; however, it is the description of the parameter that truly defines its meaning. Additionally, the sequence in which the parameters are arranged holds no significance. For each given test case, its parameter identification score is defined as follows:
$s_{PI}=s_{TS}\cdot\mathbb{I}(P=\hat{P})$ (2)
Content Filling
$s_{CF}=s_{PI}\cdot\prod_{i=1}^{N}\mathbb{I}(c_{i}=\hat{c_{i}})$ (3)
@@ 3.1 Model Selection
3.1.1 Open-Source LLMs
ToolLLaMA-2-7B-v1
ToolLLaMA-2-7B-v1 Qin et al. (2023b), developed by Tsinghua University, is a tool-oriented LLM that harnesses the power of 126,000 data samples, including more than 16,000 APIs, through supervised fine-tuning on LLaMA-2-7B-base Touvron et al. (2023b). This enables ToolLLaMA-2-7B-v1 to effectively utilize various tools to meet diverse user requirements.
ToolLLaMA-2-7B-v2
ToolLLaMA-2-7B-v2 Qin et al. (2023b) has undergone fine-tuning from LLaMA-2-7B-base, by assimilating an expansive dataset comprising over 120,000 solution paths and annotated chains of thought. To the best of our knowledge, this model stands as the most extensively trained tool-oriented LLM, utilizing the largest dataset and the broadest spectrum of tools among all available options.
NexusRaven-13B-v1
NexusRaven-13B-v1 team (2023a) is a tool-oriented model that underwent fine-tuning based on CodeLLaMA-13B Rozière et al. (2023). Distinguishing itself from prior models, NexusRaven-13B-v1 employs code nesting to invoke tools, generating the entire inference path simultaneously instead of following a step-by-step approach.
NexusRaven-13B-v2
NexusRaven-13B-v2 team (2023b) enhances the performance of NexusRaven-13B-v1 by generating single, nested, and parallel function calls in various complex scenarios. Additionally, NexusRaven-13B-v2 can generate inference paths for the function calls it creates, thereby improving overall generalization.
3.1.2 Closed-Source LLMs
GPT-3.5-turbo
GPT-3.5-turbo  [^1] https://platform.openai.com/docs/models/gpt-3-5 stands out as the most potent and cost-efficient model within the GPT-3.5 series. Tailored for conversations, it excels in comprehending and generating natural language. Furthermore, it exhibits strong tool invocation capabilities.
GPT-4
GPT-4 OpenAI (2023) represents OpenAI’s most robust LLM, surpassing its predecessor in delivering safer and more beneficial responses. Additionally, GPT-4 offers formal support for multimodal inputs and has an expanded capability to address a broader spectrum of social requirements.
3.2 Experimental Setup
In accordance with Ye et al. (2024), we adopt the ReAct Yao et al. (2023) format for inference, employing a consistent prompt template for both the ToolLLaMA-2-7B family of models and the GPT family of models. However, as NexusRaven-13B fmaily of models utilize nested functions for output, we adhere to the guidelines outlined on their official website, which necessitate the use of a distinct set of template  [^2] The specific prompt can be found in Appendix B..
3.3 Main Results
@@
As tool learning involves multiple turns of interaction between LLMs and the environment Qin et al. (2023a); Ye et al. (2024), with intricate intermediate trajectories that cannot be easily compared, our emphasis lies on evaluating the robustness of their performance in a single turn of interaction. Specifically, we evaluate the performance of various LLMs during their initial use of the tool and present the results in Table 2  [^3] The results in different scenarios can be found in Appendix E.. The resulting data reveals intriguing observations.
@@
Noise affecting tool names has a more pronounced impact on LLM performance than noise introduced to parameters. As detailed in Section 2, except for the Union-level environment, where noise is introduced to both tool names and parameters, all other noise environments introduce noise to either tool names or parameters. Consequently, we compute the absolute difference in average LLMs performance for each type of noise added to tool names or parameters, relative to their performance in the Clean-level environment, respectively. The results depicted in Figure 3 show that tool name noise significantly affects LLMs’ tool learning performance throughout the entire process. In contrast, noise in the parameters has minimal impact on the robustness of LLMs during the tool selection stage and exerts less influence on subsequent stages compared to tool name noise. Notably, LLMs exhibit greater robustness in the Union-level environment than in the Heavy (Tool) environment, underscoring the substantial impact of tool naming on model robustness.
Offering LLMs interactive examples enhances their tool learning performance, yet it does not bolster their robustness. As tool learning entails multiple turns of interaction between LLMs and external environments, we initially provide the first two turns of interactions for the test cases in each environment to evaluate LLMs’ performance during the third turn of interactions. Upon comparing GPT-4’s results in the first and third turns of interactions, as illustrated in Figure 4, it becomes evident that the provision of two turns of interaction examples leads to a consistent performance boost for GPT-4, resulting in an average performance improvement of 22.91 points across various environments. However, when examining the performance variation values, it is noteworthy that the standard deviation of its performance across environments increased from 8.14 in the first turn to 12.56 in the third turn. This observation suggests that while its performance improves, its robustness does not see a corresponding enhancement.
@@ 3.4 Why do GPT family of models NOT perform well in Slight-level environment?
A particularly intriguing finding is that, in contrast to other LLMs, the GPT family of models exhibits a lower performance in Slight-level environment compared to Medium-level, despite the limited validity of the information provided by the latter. Our thorough investigation into the model outputs has revealed that this phenomenon can be attributed to the inherent noise correction capability of the GPT family of models. For instance, when the GPT family of models selects the tool labeled as “predOict_aTge,” it automatically corrects the noise within it and generates “predict_age” as the output, consequently leading to an error   [^4] For more detailed examples, please refer to Appendix A..
@@ 4 RoTTuning
Figure 5: Illustration of RoTTuning. RoTTuning encompasses four phases (i.e., query expansion, trajectory generation, environment augmentation, and generalizability training), aiming at bolstering the robustness of LLMs in tool learning through increased environmental diversity.
In light of the experimental findings, it is evident that enhancing the robustness of LLMs in tool learning is imperative. To tackle this issue, we introduce RoTTuning, a novel approach aimed at bolstering the robustness of LLMs in tool learning through increased environmental diversity. Subsequent sections will provide a detailed explanation of this approach.
@@ 4.1 Method
As illustrated in Figure 5, RoTTuning encompasses four phases: query expansion, trajectory generation, environment augmentation, and generalizability training.
@@
To efficiently generate high-quality user queries on a large scale, we employ the self-instruct Wang et al. (2023b) technique, drawing from the 105 existing user queries  [^5] The specific prompt can be found in Appendix C.. Specifically, we instruct GPT-4 to create seven fresh user queries within the context of a subset of tools, accompanied by three existing user queries and two model-generated queries. To ensure diversity in our dataset, we scrutinize the new data for redundancy in relation to each provided example and eliminate queries with Rouge-L values surpassing 0.55. This process yields a total of 4,077 new user queries.
@@
Upon obtaining high-quality user queries, we employ GPT-4 to produce tool learning trajectories. To ensure the accuracy of the generated trajectories, we leverage the specifically designed function call feature of GPT-4. Simultaneously, we guide GPT-4 in generating the associated thought process by incorporating a system prompt  [^6] The specific prompt can be found in Appendix D.. Furthermore, we specify that GPT-4’s tool usage is limited to a maximum of nine turns. By considering each turn of interaction as a distinct data point, this process results in a total of 12,247 pieces of training data.
Environment Augmentation
All the trajectories produced by GPT-4 are initially executed in the Clean-level environment. However, to enhance the variety of environments, we aim to create trajectories that simulate noisy environments. Running GPT-4 directly in noisy environments can lead to subpar performance and compromised data quality. Therefore, we modify the trajectories generated in the Clean-level environment to align with the characteristics of noisy environments. This strategy ensures data quality while addressing the challenges of working in noisy settings. To mitigate the potential drawbacks of data coupling, we introduce randomness by augmenting 3000 trajectories for each of the Slight-, Medium-, and Heavy-level environments, along with 1500 trajectories for Union-level environments. When combined with the data from the Clean-level environment, this approach yields a total of 22,747 trajectories, representing a diverse range of environmental conditions.
Generalizability Training
Utilizing the diversity trajectories generated, we proceed with the fine-tuning of LLaMA-2-7B-base. To enhance its capability in handling scenarios where LLMs engage in multiple turns of interaction with the external environment, we implement a position interpolation Chen et al. (2023a) technique to extend its context length to 8096. Based on previous research indicating that fine-tuning with LoRA Hu et al. (2022) achieves superior generalization compared to full parametric fine-tuning Zeng et al. (2023), we opt for the LoRA fine-tuning approach. We conduct 5 epochs of training to derive the ultimate model, RoTLLaMA, which exhibits robust generalization across multiple environments.
4.2 Experimental Results
We carry out a series of experimental analyses with RoTLLaMA on RoTBench to verify its advantages when facing various noise environments.
Performance

### v3

@@ Abstract
RoTBench: A Multi-Level Benchmark for Evaluating the Robustness of Large Language Models in Tool Learning
Junjie Ye1,  Yilong Wu1,  Songyang Gao1,  Caishuang Huang1,  Sixian Li1,   Guanyu Li1,   Xiaoran Fan1,   Qi Zhang1,3thanks: Corresponding authors.,  Tao Gui2,3∗,  Xuanjing Huang1,3 1 School of Computer Science, Fudan University 2 Institute of Modern Languages and Linguistics, Fudan University 3 Shanghai Key Laboratory of Intelligent Information Processing, Fudan University jjye23@m.fudan.edu.cn {qz, tgui}@fudan.edu.cn
@@ 1 Introduction
Tool learning has emerged as a critical concept for empowering large language models (LLMs) Brown et al. (2020); Bai et al. (2022); Touvron et al. (2023a) to interact with the real world Yang et al. (2023); Mialon et al. (2023); Qin et al. (2023a); Ye et al. (2024b). In this context, the external environment of an LLM contains an ensemble of integrated tools. Each tool is uniquely identified by its name and is described by a succinct paragraph that explains its functionality. Similarly, every parameter within these tools is characterized by its name, along with a description that clarifies its purpose, its optionality, and other pertinent details.
Figure 1: Example of noise affecting tool selection for LLMs. Although the functionality of the tool remains unaffected by its name, renaming “Get_Weather” as “ABC” impedes LLMs from utilizing the tool properly.
Recent research has centered on examining how well LLMs can effectively employ tools within a carefully designed and stable environment. From one perspective, specific studies have scrutinized the outcomes of LLMs’ tool usage, verifying both the accuracy of tool selection and the efficacy of the generated responses Qin et al. (2023b); Huang et al. (2023). This analysis involved evaluating the relevance of the selected tools and the final responses in fulfilling users’ requirements. On the other hand, other investigations have delved into the intricate process of tool utilization by LLMs, striving for a more comprehensive assessment of their performance in tool learning Chen et al. (2023d); Ye et al. (2024a). This includes an analysis of the diverse capabilities necessary for LLMs to excel in tool learning while also identifying any limitations they may have in this regard.
However, these studies fail to account for the robustness of LLMs in the face of inevitable noise in real-world scenarios Chen et al. (2023b); Liu et al. (2023). Using Figure 1 as a reference, LLMs recognize the tool for querying weather information when named “Get_Weather,” but not when named “ABC,” despite the tool’s functionality remaining unaffected by its name. Consequently, it becomes imperative to investigate whether LLMs can proficiently identify these tools and configure parameters to meet user needs in noisy real-world environments. This research is essential to guarantee their reliability in practical applications.
@@
2 Related Work
Analysis of Tool Learning
Given their extensive world knowledge and superior natural language understanding, researchers have made attempts to leverage LLMs for a wide range of everyday applications Ye et al. (2023). In order to push the boundaries of their capabilities, some scholars have proposed enhancing LLMs with external tools, which has gained widespread acceptance Schick et al. (2023); Tang et al. (2023). As research in this area has deepened, certain scholars have summarized the progress made in tool learning for LLMs Mialon et al. (2023); Qin et al. (2023a), sought to uncover developmental insights, and trained more specialized LLMs for tool learning based on these findings Qin et al. (2023b); Zhuang et al. (2023); Hao et al. (2023). Furthermore, recognizing the complexity of tool learning, some researchers have specialized in evaluating not only the outcomes of tool learning Huang et al. (2023) but also the entire process Chen et al. (2023d); Ye et al. (2024a). However, it’s worth noting that all of these current efforts primarily consider LLMs’ tool usage in controlled environments, neglecting the inherent complexities of real-life scenarios. Therefore, we have undertaken an in-depth analysis of the robustness of LLMs in tool learning to advance research in a real-world context.
Robustness Testing of LLMs
Robustness is a critical factor in determining the stability of LLMs and plays a pivotal role in their practical deployment in real-life applications, which has garnered significant attention from scholars. In the early stages of research, some scholars conducted tests to assess the robustness of ChatGPT across various natural language processing tasks, highlighting the substantial room for improvement in the current robustness of LLMs Wang et al. (2023a); Chen et al. (2023c). Subsequently, other researchers specialized in creating benchmarks, such as PromptBench Zhu et al. (2023), to examine the consistency of LLM responses by introducing noise into the prompts. Given that tool learning is poised to extend the capabilities of LLMs and its outcomes can directly impact the state of the physical world Ye et al. (2024a), it becomes imperative to thoroughly evaluate its robustness.
@@ 2 Related Work
Table 1: Statistics information of the data. “# Sce”, “# Query”, “# Cat”, “# Subcat”, and “# Tool” correspond to the count of scenarios, user queries, tool categories, tool subcategories, and individual tools, respectively.
3 RoTBench
As depicted in Figure 2, RoTBench encompasses five environments, each characterized by varying levels of noise, facilitating a thorough evaluation of the robustness of LLMs throughout the three stages of tool usage.
3.1 Data Collection
In order to thoroughly cater to real-world requirements and encompass commonly utilized tools, we utilize ToolEyes Ye et al. (2024a), an evaluation system designed for tool learning. This system defines seven real-world application scenarios. Within each of these scenarios, we randomly select 15 user queries for analysis. Since the raw data offers tool information without standardized invocation paths, we have manually labeled these paths to facilitate the evaluation process. Detailed statistics of the data can be found in Table 1.
@@ 3.2 Environments Construction
To comprehensively assess the resilience of LLMs in tool learning, we reference the hierarchical classification of noise in previous studies Wang et al. (2021); Zhu et al. (2023); Dong et al. (2023) and design five distinct external environments. These environments feature varying noise levels that affect both the tool and its parameters.
@@
Slight-level environment encompasses three types of noise: insertion, omission, and substitution. These correspond to real-world occurrences such as an excess of characters, missing characters, and character errors when naming tools or parameters. Specifically, we introduce noise in the following ways: 1) We randomly select half of the available tools within the environment. For these selected tools, a random form of noise is applied, altering up to 1/3 of the characters, resulting in the creation of 105 new data points. 2) For each tool, we randomly select half of the parameters and introduce noise into their names using the method described above, generating an additional 105 new data entries. By combining these two approaches, we create a Slight-level environmental test set consisting of 210 test cases.
Medium-level environment introduces two types of noise: reversal and nonsense. These mirror real-world scenarios where names are reversed or replaced with random strings, rendering the information meaningless. To apply noise, we follow these procedures: 1) We randomly select half of the available tools. For these tools, there is a 50% probability that their names will be substituted with random strings, each containing up to 10 characters. Additionally, there is a 50% chance that the names of these tools will be reversed. This process yields 105 test cases. 2) For each tool, half of the parameters are randomly chosen. These parameters may undergo a 50% chance of having their names substituted with random strings, each containing up to 5 characters, or a 50% chance of being reversed. This leads to 105 test cases. It is worth noting that if the reversal process does not alter the name, it will be replaced with a random string. Consequently, we have successfully generated 210 test cases for the Medium-level environment.
@@
3.3 Staged Evaluation
We evaluate the robustness performance of LLMs at each of stages in tool learning and analyze their respective variations.
$s_{TS}=\mathbb{I}(t=\hat{t})$ (1)
Parameter identification involves recognizing the required parameters and outputting their respective names based on their specified needs, following the selection of the appropriate tool. This process necessitates choosing the mandatory parameters, while the optional ones are selected based on actual requirements. Similar to tool selection, the name of the parameter serves as an identifier; however, it is the description of the parameter that truly defines its meaning. For each given test case, its parameter identification score is defined as follows:
$s_{PI}=s_{TS}\cdot\mathbb{I}(P=\hat{P})$ (2)
$s_{CF}=s_{PI}\cdot\prod_{i=1}^{N}\mathbb{I}(c_{i}=\hat{c_{i}})$ (3)
@@ 4.1 Model Selection
To evaluate the robustness of widely-used LLMs with tool-use capabilities, we opt for testing four open-source models (i.e., ToolLLaMA-2-7B-v1 Qin et al. (2023b), ToolLLaMA-2-7B-v2 Qin et al. (2023b), NexusRaven-13B-v1 team (2023a), NexusRaven-13B-v2 team (2023b)) and two closed-source models (i.e., GPT-3.5-turbo [^1] https://platform.openai.com/docs/models/gpt-3-5, GPT-4 OpenAI (2023)). [^2] The details of LLMs can be found in Appendix A.
@@
4.2 Main Results
As tool learning involves multiple turns of interaction between LLMs and the environment Qin et al. (2023a); Ye et al. (2024a), with intricate intermediate trajectories that cannot be easily compared, our emphasis lies on evaluating the robustness of various LLMs during their initial use of the tool and present the results in Table 2.333The results presented are averages across various scenarios, with specific outcomes for each scenario detailed in Appendix C. The resulting data reveals intriguing observations.
@@ 4.2 Main Results
Noise affecting tool names has a more pronounced impact on LLM performance than noise introduced to parameters. We compute the absolute difference in average LLMs performance for each type of noise added to tool names or parameters, relative to their performance in the Clean-level environment, respectively. The results depicted in Figure 3 show that tool name noise significantly affects LLMs’ tool learning performance throughout the entire process. In contrast, noise in the parameters has minimal impact on the robustness of LLMs during the tool selection stage and exerts less influence on subsequent stages compared to tool name noise. Notably, LLMs exhibit greater robustness in the Union-level environment than in the Heavy (Tool) environment, underscoring the substantial impact of tool naming on model robustness.
@@
Offering LLMs interactive examples enhances their tool learning performance, yet it does not bolster their robustness. As tool learning entails multiple turns of interaction between LLMs and external environments, we initially provide the first two turns of interactions for the test cases in each environment to evaluate LLMs’ performance during the third turn of interactions. Upon comparing GPT-4’s results in the first and third turns of interactions (Figure 4), it becomes evident that the provision of two turns of interaction examples leads to a consistent performance boost for GPT-4, resulting in an average performance improvement of 22.91 points across various environments. However, when examining the performance variation values, it is noteworthy that the standard deviation of its performance across environments increased from 8.14 in the first turn to 12.56 in the third turn. This observation suggests that while its performance improves, its robustness does not see a corresponding enhancement.
@@
Figure 5: Illustration of RoTTuning. RoTTuning encompasses four phases, aiming at bolstering the robustness of LLMs in tool learning through increased environmental diversity.
A particularly intriguing finding is that, in contrast to other LLMs, the GPT family of models exhibits a lower performance in Slight-level environment compared to Medium-level, despite the limited validity of the information provided by the latter. Our thorough investigation into the model outputs has revealed that this phenomenon can be attributed to the inherent noise correction capability of the GPT family of models. For instance, when the GPT family of models selects the tool labeled as “predOict_aTge,” it automatically corrects the noise within it and generates “predict_age” as the output, consequently leading to an error.  [^4] For more detailed examples, please refer to Appendix D.
@@ 5 RoTTuning
It is evident that enhancing the robustness of LLMs in tool learning is imperative. To tackle this issue, we introduce RoTTuning, a novel approach aimed at bolstering the robustness of LLMs through increased environmental diversity.
@@ 5.1 Method
RoTTuning encompasses four phases: query expansion, trajectory generation, environment augmentation, and generalizability training (Figure 5).
@@
To efficiently generate high-quality user queries on a large scale, we employ the self-instruct Wang et al. (2023b) technique, drawing from the 105 existing user queries. [^5] The specific prompt can be found in Appendix G. Specifically, we instruct GPT-4 to create seven fresh user queries within the context of a subset of tools, accompanied by three existing user queries and two model-generated queries. To ensure diversity in our dataset, we scrutinize the new data for redundancy in relation to each provided example and eliminate queries with Rouge-L values surpassing 0.55. This process yields a total of 4,077 new user queries.
@@
Upon obtaining high-quality user queries, we employ GPT-4 to produce tool learning trajectories. To ensure the accuracy of the generated trajectories, we leverage the specifically designed function call feature of GPT-4. Simultaneously, we guide GPT-4 in generating the associated thought process by incorporating a system prompt. [^6] The specific prompt can be found in Appendix H. Furthermore, we specify that GPT-4’s tool usage is limited to a maximum of nine turns. By considering each turn of interaction as a distinct data point, this process results in a total of 12,247 pieces of training data.
@@
Environment Augmentation
To enhance the variety of environments, we modify the trajectories generated in the Clean-level environment to align with the characteristics of noisy environments. This strategy ensures data quality while addressing the challenges of working in noisy settings. To mitigate the potential drawbacks of data coupling, we introduce randomness by augmenting 3000 trajectories for each of the Slight-, Medium-, and Heavy-level environments, along with 1500 trajectories for Union-level environments. When combined with the data from the Clean-level environment, this approach yields a total of 22,747 trajectories, representing a diverse range of environmental conditions.
Generalizability Training
Utilizing the diversity trajectories generated, we proceed with the fine-tuning of LLaMA-2-7B-base Touvron et al. (2023b) and implement a position interpolation Chen et al. (2023a) technique to extend its context length to 8096. Based on previous research indicating that fine-tuning with LoRA Hu et al. (2022) achieves superior generalization compared to full parametric fine-tuning Zeng et al. (2023), we opt for the LoRA fine-tuning approach. We conduct 5 epochs of training to derive the ultimate model, RoTLLaMA, which exhibits robust generalization across multiple environments.
5.2 Experimental Results
We carry out a series of experimental analyses with RoTLLaMA on RoTBench to verify its advantages when facing various noise environments. [^7] More experiments can be found in Appendix E.
Performance
Figure 6: The means and standard deviations of our model’s performance in the five environments.